OpenStudy (anonymous):

2x^2-9x-5 one factor is -1/2 what is the other one?

OpenStudy (anonymous):

x = 5

OpenStudy (anonymous):

(2x+1)(x-5) x= -1/2 x= 5

OpenStudy (mertsj):

\[2x^2-9x-5=\frac{-1}{2}(-x^2+\frac{9}{2}x+\frac{5}{2})\]

OpenStudy (anonymous):

What is that @Mertsj ??

OpenStudy (mertsj):

Well you asked if one factor is -1/2, what is the other. If one factor of 12 is 3, what is the other?

OpenStudy (anonymous):

It's not like this @Mertsj

OpenStudy (mertsj):

So what did you really mean to ask?

OpenStudy (anonymous):

The question is this 2x^2-9x-5 one factor is -1/2 what is the other one? and the answer is so simple which is : (2x+1)(x-5) x= -1/2 x= 5

OpenStudy (mertsj):

That makes no sense. Factors are things that you multiply. If one factor is -1/2, what are you going to multiply it by?

OpenStudy (mertsj):

What you meant to say is one root is -1/2

OpenStudy (anonymous):

Let's ask @RiyaMoon and see what is the multiple choice she has?

OpenStudy (anonymous):

Hey @RiyaMoon Are you there?

OpenStudy (anonymous):

i need the other factor that goes into it. and it is 5

OpenStudy (anonymous):

thanks @RiyaMoon